### v1

@@ 4.3 A cycle with an odd number of points: An overview
Our construction uses the longest paths of Section 3.2. First we observe that our path construction can be generalized to any set of $x$-coordinates.
@@
For every even integer $n\geq 4$, every set $P$ of $n$ real numbers, and every $\delta>0$ such that the $\delta$-neighborhood of the median of $P$ does not contain any points in $P$, there exists a set $P^{\prime}$ of $n$ points in the plane with the following properties:
  1. 1.
    the $x$-projection of $P^{\prime}$ is $P$;
  2. 2.
    all $y$-coordinates are in the interval $[0,\delta]$;
  3. 3.
    the $x$-projection of any longest path on $P^{\prime}$ is a longest path on $P$;
  4. 4.
    the longest spanning path on $P^{\prime}$ is unique and noncrossing; and
  5. 5.
    the $y$-coordinates of the two endpoints of the longest path are 00 and $\delta$.
Proof sketch.
We choose the points in $P^{\prime}$ such that their $x$-coordinates are the same as the numbers in $P$ and their $y$-coordinates are in $[0,\delta]$, and thus (1) and (2) follow.
By an argument similar to the proof of Lemma 2(i) one can show that the difference of lengths of a longest and a non-longest path on $P$ is at least $2\delta$. Therefore Lemma 4 would imply that by choosing the $y$-coordinates in the interval $[0,2\delta/8k]$, any longest path on $P^{\prime}$ corresponds to a longest path on $P$, and thus (3) follows. Items (4) and (5) follow by proper choices of $y$-coordinates similar to that of Lemma 5. ∎
We can now outline the construction; see Figure 5 for an illustration. Let $n=2k{+}1$, for $k\geq 2$. We choose a set of $x$-coordinates as $P=\{-k,-(k{-}1)\epsilon,-(k{-}2)\epsilon,\ldots,-\epsilon,0,1,2,\ldots,k\}$, where $\epsilon\in(0,1/16k^{2})$ will be specified later. Note that 00 is the median of $P$, and the set $A=\{-i\cdot\epsilon:i=0,1,\ldots,k-1\}\subset[-1/16k,0]$ forms a small cluster. By Lemma 3(ii), all edges of any longest cycle on $P$ intersect the $y$-axis; and Lemma 3(iii) implies the following.
Observation 1.
The length of any cycle on $P$ that connects $p_{-k}$ to two points in $A$ is at least 2 units shorter than a longest cycle on $P$.
Figure 5: Illustration of the construction of a longest cycle for $12k{+}$ 1 points.
Below, we will specify a $y$-coordinate for each element in $P$. This will result in the point set $P^{\prime}$ for which the longest spanning cycle is unique and noncrossing. We will denote by $A^{\prime}$ the set of points in $P^{\prime}$ corresponding to $A$.
@@
It remains to specify the $y$-coordinates of the points in $P^{\prime}$ and the parameter $\epsilon$. Let $p_{x}$ denote the point in $P^{\prime}$ with $x$-coordinate $x\in P$. We first choose the $y$-coordinate for the leftmost point: Let $y_{-k}=-1/16k$; this is the only negative $y$-coordinate. We assume that $|y_{i}|\ll 1/16k$ for all other points. This ensures that the longest cycle on $P^{\prime}$ corresponds to a longest cycle on the 1-dimensional multiset where 00 represents the entire cluster $A$ (cf. Lemma 3(iii) and Lemma 7). By Lemma 3(ii), for any longest cycle on $P^{\prime}$, the two edges incident to $p_{-k}$ intersect the $y$-axis (i.e., the median). Furthermore, there is a threshold $\delta>0$ such that if $0\leq y_{i}\leq\delta$ for all remaining points, then $p_{-k}$ must be adjacent to the two closest points on or to the right of the $y$-axis: That is, $p_{-k}$ is adjacent to a point in cluster $A^{\prime}$ and to $p_{1}$ (cf. Observation 1 and Lemma 7). Next, we set $y_{0}=\delta$ and find a threshold $\delta_{1}\in(0,\delta)$ such that if $0\leq y_{i}\leq\delta_{1}$ for all remaining points and $0<\epsilon<\delta_{1}$, then the contribution of edge $p_{-k}p_{0}$ exceeds the sum of contributions of all remaining edges of a spanning cycle. Consequently, the longest cycle must include the edge $p_{-k}p_{0}$. Now both $p_{-k}$ and $p_{0}$ are fixed, and we choose a sufficiently small $\epsilon\in(0,\delta_{1})$ such that all remaining points in the cluster $A^{\prime}$ are below $p_{-k}p_{0}$ for all possible $y$-coordinates.
@@
A longest cycle on $P^{\prime}$ comprises of $p_{-k}p_{0}$, $p_{-k}p_{1}$, and the longest path $H^{\prime}$ on $P^{\prime}\setminus\{p_{-k}\}$ (from $p_{0}$ to $p_{1}$ cf. Lemma 2). By Lemma 10, we can choose $y$-coordinates for the remaining points such that $H^{\prime}$ is unique and noncrossing; and $y_{1}=0$. In particular, edge $p_{-k}p_{1}$ lies below the $x$-axis, hence below the entire path $H^{\prime}$; and $P^{\prime}\setminus\{p_{-k},p_{0}\}$ lies below the supporting line of $p_{-k}p_{0}$. Consequently, the concatenation of $p_{-k}p_{0}$, $p_{-k}p_{1}$ and $H^{\prime}$ is noncrossing.
@@ 5 Noncrossing Longest Matchings
Álvarez-Rebollar et al. [4] showed that there exist point sets for which the longest perfect matchings are noncrossing. Their example is attributed to Kåra P. Villanger in a paper by Tverberg [29]. As illustrated in Figure 6, it consists of a set $S$ of $k$ segments with endpoints in $A=\{a_{1},\dots,a_{k}\}$ and $B=\{b_{1},\dots,b_{k}\}$. The distance between any two points $a_{i}\in A$ and $b_{j}\in B$ is larger than the distance between any two points in $A$, or the distance between any two points in $B$. The points in $B$ are roughly on a vertical line. Álvarez-Rebollar et al. [4] have provided a precise description of the construction along with a detailed proof that $S$ is a longest matching for $A\cup B$.
@@
Here, we exhibit an alternative point set for which the longest perfect matching is noncrossing. Our construction follows the same framework as for paths and cycles. Let $P$ be a set of $2k$ points $p_{i}=(i,0)$ for $i=\pm 1,\pm 2,\dots,\pm k$. One can verify that a perfect matching on $P$ is longest if and only if all edges cross the $y$-axis. One such matching is $M=\{p_{-i}p_{i}\colon i=1,\dots,k\}$. Using ideas similar to those used for paths and cycles, one can assign to each $p_{i}$ a new $y$-coordinate $y_{i}$ to make $M$ longest and noncrossing at the same time; see Figure 7. The new $y$-coordinates are of the following form: $y_{1}\gg y_{-1}=y_{2}\gg y_{-2}=y_{3}\gg\cdots=y_{k}\gg y_{-k}$.
@@ 6 Some Properties of Longest Paths and Cycles
Observation 2.
Suppose that we orient the edges of a longest cycle $C$ to make it a directed cycle. Then $C$ cannot have pair of non-adjacent edges that are in convex position and have the same orientation along their convex hull.
To verify this, note that if $C$ has two such edges, say $e_{1}$ and $e_{2}$, then flipping them (replacing $e_{1}$ and $e_{2}$ by the two diagonals of the convex hull of $e_{1}$ and $e_{2}$) would produce a longer undirected cycle as in Figure 8(a). Since $e_{1}$ and $e_{2}$ have the same orientation along their convex hull, the flip does not break the cycle into two components. If every directed simple polygon $S$ contained a pair of non-adjacent edges in convex position with the same orientation along their convex hull, Observation 2 would imply Conjecture 1. However, some simple polygons do not have edges that can be flipped in this way; see e.g., Figure 8(b).
@@
To verify this observation consider an isosceles right triangle $abc$ whose right angle is at $b$. Place one point at $a$, one point at $c$, and two or more points very close to $b$. Then, the longest cycle does not contain the diametric point pair $\{a,c\}$. This observation implies that a longest cycle may not be achieved by greedily choosing longest edges.
@@
Let $S$ be a simple polygon (a noncrossing cycle) on $n$ points. Then $S$ has an edge whose length is among the smallest $83n^{2}/8+n/$ 8 distances of the $\binom{n}{2}$ point pairs.
@@
Let $e$ and $e^{\prime}$ be two edges of $S$ such that their distance along $S$ (in terms of the number of edges) is at least 2. Since $S$ is a simple polygon, $e$ and $e^{\prime}$ do not cross. Thus, there is an endpoint $p$ of $e$ and an endpoint $p^{\prime}$ of $e^{\prime}$ such that $|pp^{\prime}|$ is larger than the length of the shorter of $e$ and $e^{\prime}$, and $pp^{\prime}$ is not an edge of $S$. The number of pairs of edges at distance 00 is $n$, and the number of pairs of edges at distance $1$ is also $n$. Thus, the total number of pairs of edges at a distance at least $2$ is $\binom{n}{2}-2n$. Each such pair of edges yields a pair $\{p,p^{\prime}\}$. Each $\{p,p^{\prime}\}$ can be counted for $4$ different pairs of edges that are obtained by combining the two edges incident to $p$ and the two edges incident to $p^{\prime}$. Therefore the total number of distinct pairs $\{p,p^{\prime}\}$ is at least $\frac{1}{4}\left(\binom{n}{2}-2n\right)$. Subtracting this from the total number $\binom{n}{2}$ of point pairs yields the claimed bound. ∎
@@ Acknowledgements.
This work was initiated at the 10th Annual Workshop on Geometry and Graphs, held at Bellairs Research Institute in Barbados in February 2023. We thank the organizers and the participants.

### v2

@@ 4.3 A Cycle with an Odd Number of Points: An Overview
We start with an outline of our construction; see Figure 5 for an illustration. Let n=2k+1n=2k{+}1, for $k\geq 2$. We choose a set of $x$-coordinates as $P=\{-k,-(k{-}1)\epsilon,-(k{-}2)\epsilon,\ldots,-\epsilon,0,1,2,\ldots,k\}$, where $\epsilon\in(0,1/(16k^{2}))$ will be specified later. Note that 0 is the median of $P$, and the set $A=\{-i\cdot\epsilon:i=0,1,\ldots,k-1\}\subset[-1/(16k),0]$ forms a small cluster.
Figure 5: Illustration of the construction of a longest cycle for 2k+12k{+}1 points.
Below, we will specify a $y$-coordinate for each element in $P$. This will result in the point set $P^{\prime}$ for which the longest spanning cycle is unique and noncrossing. We will denote by $A^{\prime}$ the set of points in $P^{\prime}$ corresponding to $A$. Let $\mathcal{C}$ be the set of longest spanning cycles on $P^{\prime}$.
It remains to specify the $y$-coordinates of the points in $P^{\prime}$ and the parameter $\epsilon$. Let $p_{x}$ denote the point in $P^{\prime}$ with $x$-coordinate $x\in P$. We first choose the $y$-coordinate for the leftmost point: Let $y_{-k}=-1/(16k)$; this is the only negative $y$-coordinate. We assume that $|y_{i}|\ll 1/(16k)$ for all other points. We cannot use Corollary 1 directly for $P$ because $1/(16k)$ is too large to make any conclusion about the longest spanning cycles as the points in cluster $A^{\prime}$ are indistinguishable. Thus we define a point set, close enough to $P$, for which we can apply Corollary 1. Let $\widehat{P}$ denote the multiset obtained from $P$ by perturbing all $k$ points in $A$ to 0. Let $\mathcal{A}$ be the set of longest spanning cycles on $\widehat{P}$, and $\mathcal{B}$ the set of all other spanning cycles on $\widehat{P}$. Since the points in $\widehat{P}$ are integers, then $\beta\geq 1$, where $\beta=\min\{|A|-|B|:A\in\mathcal{A},B\in\mathcal{B}\}$. Corollary 1 implies that every longest spanning cycle $C^{\prime}$ on $P^{\prime}$ corresponds to a cycle $C\in\mathcal{A}$. This, in turn, implies that $|C^{\prime}|=|C|+Y(C^{\prime},\epsilon)$ for some small value $Y(C^{\prime},\epsilon)$ that depends on the new $y$-coordinates and $\epsilon$.
@@
In every longest cycle on $P^{\prime}$, point $p_{-k}$ is adjacent to at most one point in cluster $A^{\prime}$ and at least one point in $\{p_{1},\ldots,p_{k}\}$.
Proof.
Let $C^{\prime}$ be a longest spanning cycle on $P^{\prime}$, corresponding to a longest spanning cycle $\widehat{C}$ on $\widehat{P}$. By Lemma 5(i), every edge of $\widehat{C}$ intersects the median 0 of $\widehat{P}$. This implies that every edge of $C^{\prime}$ intersects the $y$-axis or has an endpoint in the cluster $A^{\prime}$.
Note that both $P$ and $\widehat{P}$ contain the unit-length interval $I=[0,1]$ between the leftmost k+1k{+}1 and the rightmost $k$ points. By Lemma 5(iii), n1n-1 edges of $\widehat{C}$ contain the interval $[0,1]$. This means that at most one edge of $\widehat{C}$ does not contain $[0,1]$. Consequently, in the cycle $C^{\prime}$, at most one edge connects $p_{-k}$ to a point in $A^{\prime}$ and at least one edge connects it to a point in $\{p_{1},\ldots,p_{k}\}$. ∎
@@
Similar to Lemma 11, there is a threshold $\delta>0$ such that if $0\leq y_{i}\leq\delta$ for all $i\neq-k$, then in every longest cycle on $P^{\prime}$, point $p_{-k}$ is adjacent to the two closest plausible points (subject to the constraints in Lemma 13): That is, $p_{-k}$ is adjacent to a point in cluster $A^{\prime}$ and to $p_{1}$. We set $y_{0}=\delta$ and find a threshold $\delta_{1}\in(0,\delta)$ such that if $0\leq y_{i}\leq\delta_{1}$ for all remaining points and $0<\epsilon<\delta_{1}$, then the contribution of $p_{-k}p_{0}$ and $p_{-k}p_{1}$ to $Y(C^{\prime},\epsilon)$ each exceeds the sum of contributions of all remaining edges of a cycle $C^{\prime}\in\mathcal{C}$. Consequently, every longest cycle on $P^{\prime}$ must include the edges $p_{-k}p_{0}$ and $p_{-k}p_{1}$. Now both $p_{-k}$ and $p_{0}$ are fixed, and we choose a sufficiently small $\epsilon\in(0,\delta_{1})$ such that all remaining points in the cluster $A^{\prime}$ are below $p_{-k}p_{0}$ for all possible $y$-coordinates, and that distances of the points in $A^{\prime}$ to 0 are smaller than $1/(16k)$.
Since $\epsilon>0$ is fixed now, $Y(C^{\prime},\epsilon)$ depends only on the $y$-coordinates that are still unspecified. A longest cycle $C^{\prime}$ on $P^{\prime}$ comprises of $p_{-k}p_{0}$, $p_{-k}p_{1}$, and a longest spanning path $H^{\prime}$ on $P^{\prime}\setminus\{p_{-k}\}$ from $p_{0}$ to $p_{1}$ on $P^{\prime}\setminus\{p_{-k}\}$. Let $\mathcal{H}$ denote the set of such paths $H^{\prime}$. Note that $H^{\prime}$ need not be a longest path on $P^{\prime}\setminus\{p_{-k}\}$ (in particular, $y_{0}=\delta$ may be too large to apply Corollary 1 for $P^{\prime}\setminus\{p_{-k}\}$). However, since $C^{\prime}$ corresponds to a cycle $C\in\mathcal{A}$, then all edges in $H^{\prime}$ intersect the line $y=\frac{1}{2}$ (which is the median of $P^{\prime}\setminus\{p_{-k}\}$). By Lemma 3, $H^{\prime}$ corresponds to a longest spanning path on $P\setminus\{-k\}$. As $P\setminus\{-k\}$ contains an even number of points. We can now proceed as in the proof of Lemma 6, and successively choose $y$-coordinates for the remaining points such that $H^{\prime}$ is unique and $y$-monotone (hence noncrossing); and $y_{1}=0$.
@@
Finally, we show that $C^{\prime}$ is also noncrossing. Edge $p_{-k}p_{1}$ lies below the $x$-axis, hence below the entire path $H^{\prime}$; and $P^{\prime}\setminus\{p_{-k},p_{0}\}$ lies below the supporting line of $p_{-k}p_{0}$. Consequently, the concatenation of $p_{-k}p_{0}$, $p_{-k}p_{1}$ and $H^{\prime}$ is noncrossing.
@@ 5 Noncrossing Longest Matchings
An example of a point set for which the longest perfect matching is noncrossing is already known [4]. This example is attributed to Kåra P. Villanger in a paper by Tverberg [31]. As illustrated in Figure 6, it consists of a set $S$ of $k$ segments with endpoints in $A=\{a_{1},\dots,a_{k}\}$ and $B=\{b_{1},\dots,b_{k}\}$. The distance between every two points $a_{i}\in A$ and $b_{j}\in B$ is larger than the distance between every two points in $A$, or the distance between every two points in $B$. The points in $B$ are roughly on a vertical line. A precise description of the construction along with a detailed proof that $S$ is a longest matching for $A\cup B$ is provided in [4].
@@
Here, we exhibit an alternative point set for which the longest perfect matching is noncrossing. Our construction follows the same framework as for paths and cycles. For simplicity we pick integer $x$-coordinates, however, it can be done with more general $x$-coordinates. Let $P$ be a set of $2k$ points $p_{i}=(i,0)$ for $i=\pm 1,\pm 2,\dots,\pm k$. One can verify that a perfect matching on $P$ is longest if and only if all edges cross the $y$-axis. One such matching is $M=\{p_{-i}p_{i}\colon i=1,\dots,k\}$. Using ideas similar to those used for paths and cycles, one can assign to each $p_{i}$ a new $y$-coordinate $y_{i}$ to make $M$ longest and noncrossing at the same time; see Figure 7. The new $y$-coordinates are of the following form: $y_{1}\gg y_{-1}=y_{2}\gg y_{-2}=y_{3}\gg\cdots=y_{k}\gg y_{-k}$.
@@ 6 Some Properties of Longest Paths and Cycles
Observation 1.
Suppose that we orient the edges of a longest cycle $C$ to make it a directed cycle. Then $C$ cannot have any pair of nonadjacent edges that are in convex position and have the same orientation along their convex hull.
To verify this, note that if $C$ has two such edges, say $e_{1}$ and $e_{2}$, then flipping them (replacing $e_{1}$ and $e_{2}$ by the two diagonals of the convex hull of $e_{1}$ and $e_{2}$) would produce a longer undirected cycle as in Figure 8(a). Since $e_{1}$ and $e_{2}$ have the same orientation along their convex hull, the flip does not break the cycle into two components. If every directed simple polygon $S$ contained a pair of nonadjacent edges in convex position with the same orientation along their convex hull, Observation 1 would imply Conjecture 1. However, some simple polygons do not have edges that can be flipped in this way; see for example, Figure 8(b).
@@
To verify this observation consider an isosceles right triangle abcabc whose right angle is at $b$. Place one point at $a$, one point at $c$, and two or more points very close to $b$. Then, the longest cycle does not contain the diametric point pair $\{a,c\}$. This observation implies that a longest cycle may not be achieved by greedily choosing longest edges.
@@
Let $S$ be a simple polygon (a noncrossing cycle) on $n$ points. Then the length of the shortest edge of $S$ is among the smallest $3n^{2}/8-7n/8+1$ distances of the $\binom{n}{2}$ point pairs.
@@
Let $e$ and $e^{\prime}$ be two edges of $S$ such that their distance along $S$ ( the number of edges between them) is at least 2. Since $S$ is a simple polygon, $e$ and $e^{\prime}$ do not cross. Thus, there is an endpoint $p$ of $e$ and an endpoint $p^{\prime}$ of $e^{\prime}$ such that $|pp^{\prime}|$ is larger than the length of the shorter of $e$ and $e^{\prime}$, and $pp^{\prime}$ is not an edge of $S$. The number of pairs of edges at distance 0 is $n$, and the number of pairs of edges at distance $1$ is also $n$. Thus, the total number of pairs of edges at a distance at least $2$ is $\binom{n}{2}-2n$. Each such pair of edges yields a pair $\{p,p^{\prime}\}$. Each $\{p,p^{\prime}\}$ can be counted for $4$ different pairs of edges along $S$ that are obtained by combining the two edges incident to $p$ and the two edges incident to $p^{\prime}$. Therefore, the total number of distinct pairs $\{p,p^{\prime}\}$ is at least $\frac{1}{4}\left(\binom{n}{2}-2n\right)$. Subtracting this from the total number $\binom{n}{2}$ of point pairs and then excluding the n1n-1 longest edges of $S$ yield the claimed bound for the shortest edge of $S$. ∎
@@ Acknowledgements.
This work was initiated at the 10th Annual Workshop on Geometry and Graphs, held at Bellairs Research Institute in Barbados in February 2023. We thank the organizers and the participants. We are very grateful to the anonymous reviewers who meticulously verified our proofs, and provided detailed feedback that helped to generalize main arguments in our proofs, and to improve the readability of the paper.